OpenStudy (anonymous):

Simplify the following expression: (x + 6y) - (3x − 10y).

OpenStudy (anonymous):

x+6y 3x-10y ________ now you can subtract

OpenStudy (anonymous):

subtract what?

OpenStudy (anonymous):

x + 6y- 3x + 10y = -2x +16y

OpenStudy (anonymous):

the given,,i can't just give you the answer. you must know how to solve. :-)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Let's break this down to make sure you understand the basics: x means the variable x x stands for some number Now you should also know that x means the same thing as "1 multiplied by x', because any number multiplied by 1 is just the number. So in the first parentheses above, (x + 6y), the x term just means there is one x Now 6y means "6 multiplied by y", where y is a different variable than x. It doesn't mean that y can't be the same number as x. It just means they are potentially different numbers. 6y also means six y's added together, in other words, 6y = y + y + y + y + y + y In the second parentheses, (3x - 10y), this means we have three x's and we are subtracting ten y's from the three x's. Now you also need to understand that each of the two terms in parentheses, (x + 6y), and (3x - 10y) mean the same thing as one multiplied by each term. In other words: (x + 6y) = 1(x + 6y) and (3x - 10y) = 1(3x - 10y) But also notice that we are subtracting the second parentheses term from the first one. Recall that 2 - 3 = 2 + (- 3). In other words subtracting a number means the same thing as adding the negative of the number. The same thing applies to terms in parentheses. So (x + 6y) - (3x - 10y) = (x + 6y) + [- (3x - 10y)] The terms in parentheses also represent different numbers. For instance, if x = 2 and y = 3, then (x + 6y) = (2 + 6(3)) = 20, and (3x - 10y) = (3(2) - 10(3)) = - 24 So (x + 6y) - (3x - 10y) = 20 - 24 = - 4. Now the last point to make is that - (3x - 10y) means the same thing as -1 multiplied by (3x - 10y). It also means that you are multiplying every subterm in the parentheses by -1. So - (3x - 10y) = -1(3x) - (-1)(10y) = - 3x + 10y Remember that - (- 1) = 1, the two minuses cancel each other out making a plus. This multiplying each subterm by the - 1 is an example of the "distributive property". Another example might be 5(x + 2y) = 5x + 5(2y) = 5x + 10y. In other words if there is a number outside the parentheses. Then that is the same thing as that number multiplied by each subterm inside the parentheses. That number outside the parentheses may be positive or negative. If it's negative then each subtrem is multiplied by the negative number outside the parentheses, just as we did in the -1 in our problem. In summary then, combining all this knowledge we get: (x + 6y) - (3x - 10y) = (1)(x) + (1)(6y) + (-1)(3x) - (-1)(10y) = x + 6y - 3x + 10y = x - 3x + 6y + 10y because the order of addition and subtraction doesn't make any difference = - 2x + 16y because one x minus three x's is - 2 x's, and six y's plus 10 y's is 16 y's So to answer the question what is A in the form Ax + By of our answer - 2x + 16y, of course it's - 2. this is the site might be helpful for you... http://answers.yahoo.com/question/index?qid=20120320074730AAS83o6

OpenStudy (anonymous):

Matricked,,pls.read the code of Conduct.

OpenStudy (anonymous):

Thank You So Much Trexy...! ^_^